OpenStudy (anonymous):

The water pressure during a bad neighborhood graa fire redured to 5/9 its original pressure at the hydrant. What is the reduced pressure if original pressure was 70 and 1/5 pounds per square inch?

OpenStudy (anonymous):

\[\frac{5}{9}*\left(70+\frac{1}{5}\right)=39 \text { lbs}\left/\text{\in}^2\right. \]